OpenStudy (anonymous):

r = (u+v)^2 / (x+y) solve for v

OpenStudy (anonymous):

best i got was r(x+y)= u^2 +2uv +v^2

OpenStudy (shubhamsrg):

its a quadratic eqn in v ..see ??

OpenStudy (anonymous):

\[r(x+y)=(u+v)^2\] \[\pm\sqrt{r(x+y)}=u+v\] \[v=-u\pm\sqrt{r(x+y)}\]

OpenStudy (shubhamsrg):

even without quadratic,,you can take root of both sides and solve for v .. note : you have to solve 2 times,,one with +r and other -r

OpenStudy (anonymous):

how do you know you have to take plus or minus, in satellite's answer?

OpenStudy (anonymous):

is it cause i can't realistically take the square root of a negative number? so that twill be a nonreal answer therefore being able to be discarded?

OpenStudy (shubhamsrg):

thats not unreal soln..plus and minus both come into play as both when squared yield same term.. thus his solution also gives 2 answers..which you should get from any other soln.

OpenStudy (anonymous):

my brain hurts.

OpenStudy (shubhamsrg):

lol.. ;)